# Constrained Denoising, Empirical Bayes, and Optimal Transport

### v1

@@ C.1 Variance-Constrained Denoising
  • If $G$ is sub-exponential, then we may take $\alpha=2$ and $\alpha_{n}=n^{-1}(\log n)^{c_{m}}$, for some constant $c_{m}>0$ depending only on the dimension $m$, hence $\hat{\delta}_{\mathcal{VCB}}$ achieves the rate $n^{-1}(\log n)^{c_{m}}$.
  • If $G$ is sufficiently heavy-tailed that (FM) is satisfied only for $0<\alpha<\sqrt{m^{2}+4}-m$, then we take $\alpha_{n}=n^{-(2+\alpha)(2+\alpha+m)^{-1}}(\log n)^{c_{m}}$, for some constant $c_{m}>0$ depending only on the dimension $m$, hence $\hat{\delta}_{\mathcal{VCB}}$ achieves the rate $n^{-2\alpha(2+\alpha)^{-1}}$.
@@
Second, we consider the setting of a Poisson model and nonparametric prior, where we can use either the $G$-modeling approach of the NPMLE (Example 4.1) or an $F$-modeling approach (Example 4.3) in particular the Robbins estimator Robbins [1956]. By the results of Jana et al. [2024], Polyanskiy and Wu [2021], and Shen and Wu [2022], we may derive the following rates:
  • If $G$ is compactly-supported, then we may take $\alpha=2$ and $\alpha_{n}=n^{-1}(\log n/\log\log n)^{2}$, which shows that $\hat{\delta}_{\mathcal{VCB}}$ achieves the rate of convergence $n^{-1}(\log n/\log\log n)^{2}$.
  • If $G$ is sub-exponential, then we may take $\alpha=2$ and $\alpha_{n}=n^{-1}(\log n)^{3}$, so $\hat{\delta}_{\mathcal{VCB}}$ achieves the rate of convergence $n^{-1}(\log n)^{3}$.
  • If $G$ is sufficiently heavy-tailed that (FM) is satisfied only for $20<\alpha<$ 2, then it is known that the $G$-modeling approach leads to the rate $\alpha_{n}=n^{-\alpha/(3+\alpha)}(\log n)^{c}$ for some $c>0$ and that the $F$-modeling approach leads to the rate $\alpha_{n}=n^{-(2+2\alpha)(5+2\alpha)^{-1}}(\log n)^{c^{\prime}}$ for some $c^{\prime}>0$. If $0<\alpha<\sqrt{3}-1\approx 0.7321$ then the rate of covariance estimation dominates, and both the $G$-modeling and $F$-modeling approaches lead to $\hat{\delta}_{\mathcal{VCB}}$ achieving the rate $n^{-2\alpha(2+\alpha)^{-1}}$.
@@
Last, let us consider the setting of an exponential family with its conjugate prior, as in Example 4.4. In this case, it is easy to show that the denoiser $\hat{\delta}_{\mathcal{VCB}}$ from Algorithm 1 has the following formula:
@@
where $\hat{\mu}:=\frac{1}{n}\sum_{i=1}Z_{i}$ and $\hat{\sigma}^{2}:=\frac{1}{n-1}\sum_{i=1}^{n}(Z_{i}-\hat{\mu})^{2}$ are the empirical mean and variance of $Z_{1},\ldots,Z_{n}$. Since we can take $\alpha_{n}=n^{-1}$ and $\alpha=2$, we obtain the expected behavior, that $\hat{\delta}_{\mathcal{VCB}}$ achieves the parametric convergence rate $n^{-1}$.
C.2 Distribution-Constrained Denoising
@@
First, consider the setting of a Gaussian model and nonparametric compactly supported latent variable distribution $G$. We already saw in Appendix C that assumptions (2M), (ZC), and (BC) are satisfied, and that Soloff et al. [2024, Theorem 9] allows us to take $\alpha_{n}=n^{-1}(\log n)^{c_{m}}$ for some constant $c_{m}>0$ depending only on the dimension $m$. We may also use Soloff et al. [2024, Theorem 11] to take $\beta_{n}=(\log n)^{-1}$ which is known to be the minimax optimal rate for Gaussian deconvolution. (See Dedecker and Michel [2013, Theorem 1].) Lastly, we use Chewi et al. [2025, equation (2.21)] to take
$\gamma_{n}:=\begin{cases}n^{-\nicefrac{{1}}{{2}}}&\textnormal{ if }1\leq m\leq 3\\ n^{-\nicefrac{{1}}{{2}}}\log n&\textnormal{ if }m=4\\ n^{-\nicefrac{{2}}{{m}}}&\textnormal{ if }m\geq 5.\end{cases}$ (C.1)
@@
Second, consider the setting of an exponential family with its conjugate prior, as in Example 4.4. As discussed in Appendix C, assumptions (2M), (ZC), and (BC) are satisfied, and we may take $\alpha_{n}=n^{-1}$. Also, it is expected that one can typically take $\beta_{n}=n^{-1}$. Lastly, note that we may also apply [Chewi et al., 2025, equation (2.21)] to determine that $\gamma_{n}$ is given by (C.1). For any dimension $m$, the empirical Wasserstein distance dominates (since it is always at least as slow as $n^{-\nicefrac{{1}}{{2}}}$), and we conclude that $\hat{\delta}_{\mathcal{DCB}}$ achieves the rate $\gamma_{n}$.
Appendix D Proofs from Section 3

### v2

@@ B.1 Variance-Constrained Denoising
  • If $G$ is sub-exponential, then we may take $q=4$ and $\alpha_{n}=n^{-1}(\log n)^{c_{m}}$, for some constant $c_{m}>0$ depending only on the dimension $m$, hence $\hat{\delta}_{\mathcal{VCB}}$ achieves the rate $n^{-1}(\log n)^{c_{m}}$.
  • If $G$ is sufficiently heavy-tailed that (FM) is satisfied only for $2<q<2+\sqrt{m^{2}+4}-m$, then we take $\alpha_{n}=n^{-q(q+m)^{-1}}(\log n)^{c_{m}}$, for some constant $c_{m}>0$ depending only on the dimension $m$, hence $\hat{\delta}_{\mathcal{VCB}}$ achieves the rate $n^{-2(q-2)/q}$.
@@
Second, we consider the setting of a Poisson model and nonparametric prior, where we can use either the $G$-modeling approach of the NPMLE (Example 4.1) or an $F$-modeling approach (Appendix A.1) in particular the Robbins estimator Robbins [1956]. By the results of Jana et al. [2024], Polyanskiy and Wu [2021], and Shen and Wu [2022], we may derive the following rates:
  • If $G$ is compactly-supported, then we may take $q=4$ and $\alpha_{n}=n^{-1}(\log n/\log\log n)^{2}$, which shows that $\hat{\delta}_{\mathcal{VCB}}$ achieves the rate of convergence $n^{-1}(\log n/\log\log n)^{2}$.
  • If $G$ is sub-exponential, then we may take $q=4$ and $\alpha_{n}=n^{-1}(\log n)^{3}$, so $\hat{\delta}_{\mathcal{VCB}}$ achieves the rate of convergence $n^{-1}(\log n)^{3}$.
  • If $G$ is sufficiently heavy-tailed that (FM) is satisfied only for $42<q<$ 4, then it is known that the $G$-modeling approach leads to the rate $\alpha_{n}=n^{-(q-2)(q+1)^{-1}}(\log n)^{c}$ for some $c>0$ and that the $F$-modeling approach leads to the rate $\alpha_{n}=n^{-(2q-2)(2q+1)^{-1}}(\log n)^{c^{\prime}}$ for some $c^{\prime}>0$. If $2<q<\sqrt{3}+1\approx 2.7321$ then the rate of covariance estimation dominates, and both the $G$-modeling and $F$-modeling approaches lead to $\hat{\delta}_{\mathcal{VCB}}$ achieving the rate $n^{-2(q-2)/q}$.
@@
Last, let us consider the setting of an exponential family with its conjugate prior, as in Appendix A.2. As we already discussed, the denoiser $\hat{\delta}_{\mathcal{VCB}}$ from Algorithm 1 has the following formula:
@@
where $\hat{\mu}:=\frac{1}{n}\sum_{i=1}Z_{i}$ and $\hat{\sigma}^{2}:=\frac{1}{n-1}\sum_{i=1}^{n}(Z_{i}-\hat{\mu})^{2}$ are the empirical mean and variance of $Z_{1},\ldots,Z_{n}$. Since we can take $\alpha_{n}=n^{-1}$ and $q=4$, we obtain the expected behavior, that $\hat{\delta}_{\mathcal{VCB}}$ achieves the parametric convergence rate $n^{-1}$.
B.2 Distribution-Constrained Denoising
@@
First, consider the setting of a Gaussian model and nonparametric compactly supported latent variable distribution $G$. We already saw in Appendix B that assumptions (2M), (ZC), and (BC) are satisfied, and that Soloff et al. [2024, Theorem 9] allows us to take $\alpha_{n}=n^{-1}(\log n)^{c_{m}}$ for some constant $c_{m}>0$ depending only on the dimension $m$. We may also use Soloff et al. [2024, Theorem 11] to take $\beta_{n}=(\log n)^{-1}$ which is known to be the minimax optimal rate for Gaussian deconvolution. (See Dedecker and Michel [2013, Theorem 1].) Lastly, we use Chewi et al. [2025, equation (2.21)] to take
$\gamma_{n}:=\begin{cases}n^{-\nicefrac{{1}}{{2}}}&\textnormal{ if }1\leq m\leq 3\\ n^{-\nicefrac{{1}}{{2}}}\log n&\textnormal{ if }m=4\\ n^{-\nicefrac{{2}}{{m}}}&\textnormal{ if }m\geq 5.\end{cases}$ (B.1)
@@
Second, consider the setting of an exponential family with its conjugate prior, as in Subsection A.2. As discussed in Appendix B, assumptions (2M), (ZC), and (BC) are satisfied, and we may take $\alpha_{n}=n^{-1}$. Also, it is expected that one can typically take $\beta_{n}=n^{-1}$. Lastly, note that we may also apply [Chewi et al., 2025, equation (2.21)] to determine that $\gamma_{n}$ is given by (B.1). For any dimension $m$, the empirical Wasserstein distance dominates (since it is always at least as slow as $n^{-\nicefrac{{1}}{{2}}}$), and we conclude that $\hat{\delta}_{\mathcal{DCB}}$ achieves the rate $\gamma_{n}$.
Appendix C Results on General-Constrained Denoising
In this section we detail the oracle- and empirical-level versions of problem ($\mathcal{GCB}$), which were mentioned but not precisely studied in the main body of the paper.
C.1 Oracle General-Constrained Denoising
We begin with the oracle version of problem ($\mathcal{GCB}$) of general-constrained Bayes denoising. That is, we fix a linearly-independent collection $\psi_{1},\ldots,\psi_{k}:\mathbb{R}^{m}\to\mathbb{R}$ of continuous functions (which are also assumed to be $G$-integrable). For convenience, we write $\Psi:=\{\psi_{1},\ldots,\psi_{k}\}$. Roughly speaking, any constraint on the distribution of $\delta(Z)$ can be cast in the form above (although we may need to take $k$ to be infinite); indeed, ($\mathcal{GCB}$) reduces to ($\mathcal{B}$) if $\Psi$ is empty, ($\mathcal{VCB}$) if $\Psi$ consists of all quadratic forms, and ($\mathcal{DCB}$) if $\Psi$ consists of all bounded continuous functions. In addition to subsuming these earlier examples, we are motivated by the following.
Example C.1 (Higher moment constraints).
$\begin{cases}\textnormal{minimize}&\mathbb{E}\left[|\delta(Z)-\Theta|^{2}\right]\\ \textnormal{over}&\delta:\mathbb{R}\to\mathbb{R}\\ \textnormal{with}&\mathbb{E}\left[(\delta(Z))^{\ell}\right]=\mathbb{E}\left[\Theta^{\ell}\right]\quad\textnormal{ for all }\ell=1,\ldots,k.\end{cases}$ (C.1)
This, of course, coincides with the case of variance constraints when $k=2$, and higher moment constraints force the distribution of $\delta(Z)$ to be even closer to the distribution $G$ of $\Theta$. Along these lines it is shown in Armstrong et al. [2022], in the context of constructing nonparametric EB confidence intervals, that imposing constraints on the fourth moment leads to EB procedures which nearly adapt to the parametric (Gaussian) setting.
Example C.2 (Errors-in-variables regression continued: moment adjusted imputation).
Bay [1997] and Thomas et al. [2011] propose an intermediate approach called moment adjusted imputation that bridges the methods described in Examples 3.2 and 3.7. Their approach uses a denoiser that matches a controlled number of moments and cross-moments between $X$ and $Y$, extending beyond the two moments in moment reconstruction but without enforcing full distributional equivalence. For instance, when $X\in\mathbb{R}$, they aim to estimate $X$ while preserving multiple moments such as $\mathbb{E}[X]$, $\mathbb{E}[X^{2}]$, $\mathbb{E}[X^{3}]$, $\mathbb{E}[X^{4}]$, as well as cross-moments like $\mathbb{E}[XY]$ and $\mathbb{E}[X^{2}Y]$. Such a denoiser is given by problem ($\mathcal{GCB}$) with the choices $Z=(W,Y)$, $\Theta=(X,Y)$, $\psi_{\ell}(x,y)=x^{\ell}$ for $\ell=1,\dotsc,4$, and $\psi_{5}(x,y)=xy$, $\psi_{6}(x,y)=x^{2}y$.
Furthermore, we assume that
[…]
We note that (C) is a mild condition, and that it is satisfied whenever $\{P_{\theta}\}_{\theta}$ is a suitable exponential family, regardless of $G$ [García-Trillos and Sen, 2024, Remark 2.2]; for example, it holds for Gaussian or Poisson families. (It holds trivially whenever $F$ has discrete support.) Lastly, we assume that
[…]
Theorem C.3.
Under assumptions (ZC), (BC), (C), and (QG), the problem
$\underset{\pi\in\Gamma(F;G,\Psi)}{\textnormal{minimize}}\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}(z,\eta)\,\textnormal{d}\pi(z,\eta),$ (C.2)
has a unique solution $\pi_{\mathcal{GCB}}$, and this solution is concentrated on the graph of a function which we denote $\delta_{\mathcal{GCB}}:\mathbb{R}^{d}\to\mathbb{R}^{m}$. If $H$ denotes the marginal distribution of $D$ when $(Z,D)\sim\pi_{\mathcal{GCB}}$, then $\delta_{\mathcal{GCB}}$ can be written as $\delta_{\mathcal{GCB}}=\nabla\phi\circ\delta_{\mathcal{B},H}$ for some convex function $\phi:\mathbb{R}^{m}\to\mathbb{R}\cup\{+\infty\}$, and we have $R_{\mathcal{GCB}}=R_{\mathcal{B}}+W_{2}^{2}(H,(\delta_{\mathcal{B}})_{\#}F)$. Consequently, (C.2) is a tight convex relaxation of ($\mathcal{GCB}$), and $\delta_{\mathcal{GCB}}$ is the $F$-a.s. unique solution to ($\mathcal{GCB}$).
As in the case of problem ($\mathcal{DCB}$), this result provides conditions under which problem ($\mathcal{GCB}$) has a unique solution which can be found by solving the linear program (C.2). We refer to the probability measure as $H$ as the projected latent variable distribution since, among all probability measures whose integrals with $\Psi$ match those of $G$, it is the nearest to $(\delta_{\mathcal{B}})_{\#}F$ with respect to the distance $W_{2}$.
Example C.4 (Higher moment constraints continued).
To provide further intuition on the projected latent variable distribution $H$, we characterize it in some particular cases of Example C.1. If $k=0$ or $k=1$, then $\delta_{\mathcal{B}}$ is feasible hence optimal for (C.1), so $H$ equals $(\delta_{\mathcal{B}})_{\#}F$. If $k=2$, then (C.1) is equivalent to ($\mathcal{VCB}$), so $H$ equals the pushforward of $F$ by an affine function of $\delta_{\mathcal{B}}$. If $G$ is compactly supported and $H_{k}$ denotes the projected latent variable distribution for constraints on $k$ moments, then $H_{k}$ converges weakly to $G$ as $k\to\infty$.
C.2 Empirical General-Constrained Denoising
Since in Theorem C.3 we provided sufficient conditions for the existence of a unique solution to problem ($\mathcal{GCB}$), we presently develop an EB methodology for targeting this oracle denoiser. Methodologically, we regard ($\mathcal{GCB}$) as a way to tradeoff between problems ($\mathcal{VCB}$) and ($\mathcal{DCB}$); the former only encodes limited information about the latent variable distribution (the first two moments), the latter suffers from slow rates of convergence (due to the difficulty of the deconvolution problem) and may have larger denoising risk. Often some intermediate behavior may be desired.
We describe our proposed procedure in Algorithm 3, which is similar to the approach we previously used for the distribution-constrained problem ($\mathcal{DCB}$). Since we can estimate the marginal distribution $F$ via $\bar{F}_{n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{Z_{i}}$ and the cost function $c_{G}$ via $\hat{c}_{n}$ through $\hat{\delta}_{\mathcal{B}}$, it only remains to approximate the constraints of the linear program (C.2). Assuming that this is possible (that is, that the integrals of $\Psi$ with $\hat{G}_{n}$ converge to the integrals of $\Psi$ with $G$), we may let $\hat{\delta}_{\mathcal{GCB}}$ denote the barycentric projection of an optimal solution to the corresponding linear program. Note also that we restrict the $\eta$ variable of the linear program to lie in a compact set $K\subseteq\mathbb{R}^{m}$ (line 6).
In order to give a rate of convergence of $\hat{\delta}_{\mathcal{GCB}}$ to $\delta_{\mathcal{GCB}}$, we need to present a set of assumptions. To set this up, we first assume (ZC), (BC), (C), and (QG) so that Theorem C.3 is in effect; then, we recall that there is a well-defined projected latent variable distribution which we denote $H$. Second, we assume
in probability as $n\to\infty$. For $\psi_{\ell}$ with at most quadratic growth, such convergence follows from (EC) because of Villani [2009, Definition 6.8]. For $\psi_{\ell}$ whose growth is polynomial of degree $p\geq 2$, it is sufficient to additionally require $W_{p}(\hat{G}_{n},G)=o_{\mathbb{P}}(1)$ on top of (EC).
$W_{2}^{2}(\hat{H}_{n},H)=O_{\mathbb{P}}(\delta_{n}).$ (D)
which is analogous to (EW).
5:     $\hat{G}_{n}\leftarrow$ estimate of $G$
8:                 over         probability measures $\pi\in\mathcal{P}(\mathbb{R}^{d}\times K)$
9:                 with        $\pi(\{Z_{i}\}\times K)=\frac{1}{n}$ for all $1\leq i\leq n$
12:     return $\hat{\delta}_{\mathcal{GCB}}$
13:end procedure
Finally, these considerations lead us to the following result.
Theorem C.5.
$\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{GCB}}(Z_{i})-\delta_{\mathcal{GCB}}(Z_{i})\right\|^{2}=O_{\mathbb{P}}(\alpha_{n}^{\nicefrac{{1}}{{2}}}\vee\delta_{n}\vee\varepsilon_{n})$ (C.3)
as $n\to\infty$.
The conclusion of Theorem C.5 for the denoiser $\hat{\delta}_{\mathcal{GCB}}$ given in Algorithm 3 is that the rate of convergence of general-constrained Bayes denoising is determined by whichever of unconstrained Bayes denoising, convergence of the projected latent variable distributions, and convergence of the empirical Wasserstein distance is slowest. The rate of convergence of the projected latent variable distributions, described by $\delta_{n}$, exactly quantifies the statistical tradeoff of interpolating between problems ($\mathcal{VCB}$) and ($\mathcal{DCB}$). In the cases we consider in Example C.4, we may anticipate $\delta_{n}$ to decay at nearly parametric rates when $k=0$ (no constraints), or when $k=2$ (variance constraints) and the empirical Bayes denoiser may be estimated at fast rates (as discussed in Example 4.1), or inversely polylogarithmic in $n$ when $k\to\infty$ (and we effectively need to deconvolve the latent variable distribution).
Appendix D Details for Heterogeneous (Empirical) Constrained Bayes
The unconstrained denoising problem in model (5.1) is as follows:
$\begin{cases}\textnormal{minimize}&\mathbb{E}\left[\|\delta(Z;\Xi)-\Theta\|^{2}\right]\\ \textnormal{over}&\delta:\mathbb{R}^{d}\times\mathbb{R}^{p}\to\mathbb{R}^{m}.\end{cases}$ (D.1)
$\delta_{\mathcal{B}}(z;\xi)=\mathbb{E}[\Theta\,|Z=z,\Xi=\xi].$
In the next subsections, we consider adding constraints to problem (D.1). While we will not prove rigorous results regarding asymptotic theory, we extend the strategies of the previous sections to accommodate the heterogeneity. An important feature of heterogeneous constrained denoising is that we now must distinguish between two kinds of constraints: the so-called marginal constraints on the distribution of $\delta_{\mathcal{B}}(Z;\Xi)$, and the so-called conditional constraints on the distribution of $\delta_{\mathcal{B}}(Z;\Xi)$ given $\{\Xi=\xi\}$.
Before explaining how to extend our constrained denoising methods to model (5.1), we first discuss possible applications of the model in more detail.
D.1 Applications of the heterogeneous setting
Model (5.1), allows for heterogeneity in both the prior $(G_{\xi})$ and the likelihood $P_{\theta,\xi}$. Two important examples for heterogeneity in the likelihood are the following:
  • (Heteroskedastic Gaussian.) For $m=d$, assume that the parameter $\theta$ is a vector in $\mathbb{R}^{m}$ and the parameter $\xi$ is a positive semi-definite $m\times m$ matrix, which we choose to denote by $\Sigma$. Let $P_{\theta,\Sigma}=\mathcal{N}(\theta,\Sigma)$. As one concrete application, suppose that conditional on $\Theta_{i}$, $Z_{i}$ is a sample average of $K_{i}$ i.i.d. observations $Z_{ij}$ with $Z_{ij}\sim\mathcal{N}(\Theta_{i},\sigma^{2})$. Then $Z_{i}\sim\mathcal{N}(\Theta_{i},\sigma_{i}^{2})$ with $\sigma_{i}^{2}=\sigma^{2}/K_{i}$. The heteroskedastic Gaussian setting encompasses this situation by taking $m=d=1$ and $\Xi_{i}=\sigma^{2}/K_{i}$.
  • (Poisson with heterogeneous exposure.) For $m=d=1$, the parameter $\theta$ is a non-negative real number and the parameter $\xi$ is a non-negative real number, which we choose to denote by $\lambda$. Then let $P_{\theta,\lambda}=\textnormal{Poi}(\lambda\theta)$. We call $\lambda$ the exposure [Norberg, 1979]. The intuition here is that we are interested in the mean of the $i$-th observation normalized by the exposure time, that is, $\Theta_{i}=\mathbb{E}[Z_{i}/\lambda_{i}]$; see Subsection 6.2 for an application to baseball in which $\lambda_{i}$ is the number of games played by the $i$-th player.
Moreover, in several applications, the distribution of $\Theta$ may depend on $\Xi$. Such dependence implies that $\Theta_{1},\ldots,\Theta_{n}$ are no longer exchangeable. Examples where this situation occurs are as follows.
  • (Heteroskedastic Gaussian, continued.) In our application to astronomy (Subsection 6.1), we assume that $\Theta$ is independent of $\Xi\equiv\Sigma$, and as mentioned therein, this is a common assumption in the literature. However, several authors have noted that this assumption can be violated [Stephens, 2017, Gu and Koenker, 2017b, Weinstein et al., 2018, Chen, 2022, Ignatiadis and Sen, 2025]. If we model $G_{\xi}$ as a function of $\xi$ (using, e.g., the G-modeling approach of Chen [2022]), then we can address this issue (and the general model in (5.1) accommodates this dependence).
  • (Side-information.) In several common applications, we observe covariates (features) $X_{i}$ in addition to the noisy observations $Z_{i}$ for each $i$. The covariates $X_{i}$ are informative about $\Theta_{i}$, though commonly they do not directly impact the likelihood, that is, $Z_{i}$ is independent of $X_{i}$ conditionally on $\Theta_{i}$. The setting with side-information falls under our framework by letting $\Xi_{i}=X_{i}$ and $P_{\theta,\xi}=P_{\theta}$. Examples of empirical Bayes modeling with side-information include Fay III and Herriot [1979], Cohen et al. [2013], Ignatiadis and Wager [2019], Banerjee et al. [2023].
  • (Heteroskedasticity and side-information.) It is possible to have both side-information and heterogeneity in the likelihood at the same time, such as described in Chen [2022], Kwon [2023], Ghosh et al. [2025]. We can model this situation, e.g., in the heteroskedastic Gaussian setting, by letting $\Xi_{i}=(\Sigma_{i},X_{i})$. Herein, $P_{\theta,\xi}$ only depends on the likelihood covariance (not on the side-information), while the prior can depend on both $\Sigma_{i}$ and $X_{i}$.
D.2 Variance-Constrained Denoising
D.2.1 Marginal Variance-Constrained Denoising
For the first approach, we require that the marginal mean and variance of the denoised data match the marginal mean and variance of the latent variable. That is, we solve problem (D.1) subject to the constraints
$\mathbb{E}[\delta(Z;\Xi)]=\mathbb{E}[\Theta]\qquad\textnormal{ and }\qquad\textnormal{Cov}(\delta(Z;\Xi))=\textnormal{Cov}(\Theta).$ (D.2)
This approach was also followed by Ghosh [1992] in a small area application with heterogeneous sample sizes across areas. By noticing that this is closely related to the setting of Section 3.1 but with observation $\tilde{Z}:=(Z,\Xi)$, it can be shown that the oracle denoiser is
$\hat{\mu}:=\int_{\mathbb{R}^{m}}\theta\textnormal{d}\hat{G}_{n}(\theta)\qquad\text{ and}\qquad\,\hat{A}:=\int_{\mathbb{R}^{m}}(\theta-\hat{\mu})(\theta-\hat{\mu})^{\top}\textnormal{d}\hat{G}_{n}(\theta).$ (D.3)
D.2.2 Conditional Variance-Constrained Denoising
Another possible constraint is that the conditional mean and covariance of $\delta(Z;\Xi)$ match the conditional mean and covariance of $\Theta$ given $\Xi$. In other words, we may solve problem (D.1) subject to the constraints
$\mathbb{E}[\delta(Z;\Xi)\,|\,\Xi]=\mathbb{E}[\Theta\,|\,\Xi]\qquad\textnormal{and }\qquad\textnormal{Cov}(\delta(Z;\Xi)\,|\,\Xi)=\textnormal{Cov}(\Theta\,|\,\Xi)\qquad\textnormal{ almost surely.}$. (D.4)
In this setting, it can be shown that the oracle denoiser is
where $\mu=\mathbb{E}[\Theta]$. Our EB approach will target the denoiser $\delta_{\mathcal{C}\mathcal{VCB}}$ by plugging in suitable estimates for $\mathbb{E}[\Theta],\textnormal{Cov}(\Theta\,|\,\Xi=\xi)$, and $\textnormal{Cov}(\delta_{\mathcal{B}}(Z,\Xi)\,|\,\Xi=\xi)$. If we additionally assume $\Theta\perp\!\!\!\!\perp\Xi$ (as we will assume in our later applications), then we may estimate $\mathbb{E}[\Theta]$ and $\textnormal{Cov}(\Theta)\equiv\textnormal{Cov}(\Theta\,|\,\Xi)$ in the same way as above, via the corresponding moments of the NPMLE given in (D.3). However, $\textnormal{Cov}(\delta_{\mathcal{B}}(Z,\Xi)\,|\,\Xi=\xi)$ must be handled more carefully. If we are in the setting of $G$-modeling and the measure $\hat{G}_{n}$ is finitely supported (which is true for the NPMLE), then we may approximate
via
$\sum_{j=1}^{r}w_{j}\int_{\mathbb{R}^{d}}(\hat{\delta}_{\mathcal{B}}(z,\xi)-\mu)(\hat{\delta}_{\mathcal{B}}(z,\xi)-\mu)^{\top}P_{\theta_{j},\xi}(\textnormal{d}z).$ (D.5)
D.2.3 Numerical Illustration
In Figure D.1 we show the result of applying three denoisers to this data set. The marginal variance-constrained denoiser (third column) leads to the incorrect variance when applied only to the low-variance observations (second row) or the high-variance observations (third row). Contrarily, the conditional variance-constrained denoiser (fourth column) obtains the correct variance for both subsets of the data.
D.3 Marginal Distribution-Constrained Denoising
Next we seek to extend the distribution-constrained problem ($\mathcal{DCB}$) to the heterogeneous setting. To state the problem precisely, we aim to solve problem (D.1) subject to the constraint that the marginal distribution of the denoised data matches the marginal distribution of the latent variable, meaning
$\delta(Z;\Xi)\overset{\mathcal{D}}{=}\Theta.$ (D.6)
and we replace barycentric projection (line 12) by
Figure D.1: Denoising in the heterogeneous setting of Section D to a toy data set. By rows, we show histograms of all of the observations (first row), the subset of low-variance observations (second row), and the subset of high-variance observations (third row). By columns, we show the raw data (first column), the EB denoised data set (second column), the marginal variance-constrained EB denoised data set (third column), and the conditional variance-constrained EB denoised data set (fourth column). In each plot we show the density of the latent variable distribution (blue) for convenience.
D.4 Marginal General-Constrained Denoising
$\mathbb{E}[\psi_{\ell}(\delta(Z,\Xi))]=\mathbb{E}[\psi_{\ell}(\Theta)]\quad\textnormal{ for all }1\leq\ell\leq k,$ (D.7)
where $\psi_{1},\ldots,\psi_{k}$ are some fixed functions. As in Subsection 3.3, the oracle denoiser in this problem is the composition of the Bayes denoiser $\delta_{\mathcal{B}}$ with an OT map from the distribution of $\delta_{\mathcal{B}}(Z,\Xi)$ to some distribution $H$ (whose integrals with respect to $\Psi$ agree with those of $G$) but which is not a priori known. Nonetheless, this denoiser can be approximated via applying the same modifications of Subsection D.3 to Algorithm 3.
Appendix E Computational Considerations
In this section, we briefly describe some computational considerations that arise in the implementation of various algorithms described in the main body of the paper. Our procedures also inherit some computational complexity from the unconstrained EB denoisers which they take as input. Standard OT problems are solved with the Python Optimal Transport (POT) package [Flamary et al., 2021], and other convex optimization problems are solved via CVXPY [Diamond and Boyd, 2016] using the Mosek  [ApS, 2025] interior point convex programming solver.
E.1 $G$-modeling via nonparametric maximum likelihood
  1. 1.
  2. 2.
In the Gaussian models (Figure 1.1, Figure D.1, and Figure 6.1) we inialize the NPMLE atoms in Step 1 to be the data points themselves, following the so-called exemplar$+$ method of Soloff et al. [2024]. In the Poisson model of Figure 6.2 we initialize the NPMLE atoms of step 1 to be a $50\times 50$ equispaced grid for the minimal bounding box of the data.
E.2 Variance-Constrained Denoising
Second, we consider the two variance-constrained denoising problems in the heterogeneous setting of Section D. The marginal variance-constrained denoiser $\hat{\delta}_{\mathcal{M}\mathcal{VCB}}$ does not require any further computational burden compared to $\hat{\delta}_{\mathcal{VCB}}$, at least in the case where $\hat{\delta}_{\mathcal{B}}$ arises from $G$-modeling. This is the case in all of our applications except for Figure D.1. The conditional variance-constrained denoiser $\hat{\delta}_{\mathcal{C}\mathcal{VCB}}$ is more complicated since it requires the numerical integration of the terms in (D.5) for each evaluation of $\hat{\delta}_{\mathcal{C}\mathcal{VCB}}(Z_{i},\Xi_{i})$; we compute these integrals via Monte Carlo with 100 trials each.
E.3 Distribution- and General-Constrained Denoising
Our procedures Algorithm 2 and Algorithm 3 for approximating the solution to the distribution-constrained problem ($\mathcal{DCB}$) and the general-constrained problem ($\mathcal{GCB}$) both require numerically solving a linear program (lines 7 to 10), so we briefly discuss some details of this. For problem ($\mathcal{DCB}$) we use the solver from POT which is specifically designed to handle OT problems, and for problem ($\mathcal{GCB}$) we use Mosek’s general-purpose solvers for linear programming in CVXPY. In either case, some discretization of the $\eta\in\mathbb{R}^{m}$ coordinate is necessary. We always choose the discretization to be an equispaced grid on some axis-aligned bounding box, and the main difficulty is choosing an appropriate discretization.
In the cases where $m=1$ (i.e., Figure D.1), the complexity of the linear program problem scales linearly in the number of discretization points $k$. So, it is possible to take $k$ quite large without computational burden; for the sake of simplicity. In these cases, we fixed the discretization of $\eta$ to consist of $k=200$ equispaced points in the minimal bounding interval of the observations.
In the cases where $m\geq 2$ (i.e., Figure 1.1, Figure 6.2, and Figure 6.1) the computation is more burdensome. We take $k=300$ in problem ($\mathcal{DCB}$) since the POT package is highly optimized for this task, and we take $k=200$ for problem ($\mathcal{GCB}$) although computation is a bit slow.
Additionally, we remark that naively discretizing the linear program in Algorithm 3 may lead to infeasibility; this is because $\hat{G}_{n}$ witnesses feasibility for arbitrary probability measures but not necessarily for probability measures supported on the prescribed grid points. To avoid this, we augment the discretization by adding the NPMLE atoms in addition to the grid points.
Appendix F Proofs from Section 3